# Kennedy-Tasaki transformation and non-invertible symmetry in lattice models beyond one dimension

### v1

@@ Acknowledgements.
The authors would like to thank Shu-Heng Shao and Yunqin Zheng for useful discussions. APM and HS would like to thank Takuya Okuda for discussions on a project that uses similar ideas. This work was supported by the National Science Foundation under Award No. PHY 2310614. T.-C.W. also acknowledges the support by Stony Brook University’s Center for Distributed Quantum Processing.
@@ A.1 Majorana plaquette model
Hence, in total, there are $2^{L_{x}+L_{y}-1}$ subsystem symmetries, which in turn agrees with the number of subsystem symmetries for the plaquette Ising model. Now we discuss gauging the subsystem fermion parity. We consider the Jordan-Wigner transformation for subsystem symmetric fermionic models in 2D defined in Ref. [29],
@@
$\begin{split}H_{plaq}&=-\sum_{i\neq L,j}Z_{i,j}Z_{i+1,j}Z_{i,j+1}Z_{i+1,j+1}-\sum_{j=1}^{L_{y}}\left(\prod_{i=1}^{L_{x}}X_{i,j}\right)\left(\prod_{i=1}^{L_{x}}X_{i,j+1}\right)Z_{L_{x},j}Z_{1,j}Z_{L_{x},j+1}Z_{1,j+1}-\lambda\sum_{i,j}X_{i,j}.\end{split}$ (126)
This is the plaquette Ising model Hamiltonian with defects along the horizontal direction. The second term in Eq. (126) contain defects inserted at two consecutive horizontal lines. Note that the subsystem fermion parity maps to the subsystem line symmetry of the plaquette Ising model under the Jordan-Wigner transformation Eq. (A.1),
@@
We perform a procedure similar to the procedure of gauging fermion parity in the free Majorana fermion model to obtain the transverse field Ising model in 1D [16]. First let us define an extended Hilbert space.
$\mathcal{H}\equiv\mathcal{H}_{0}\bigoplus_{\{i_{1},...,i_{k}\}\neq\phi}\mathcal{H}_{i_{1}...i_{k}},\qquad\{i_{1},...,i_{k}\}\subset\{1,...,L_{y}\}$ (128)
where $\phi$ is the empty set and $i_{1}<...<i_{k}$. Now let us introduce a total ordering on the subsets of the set $\{1,2,...,L_{y}\}$. Suppose $A$ and $B$ are two subsets of $\{1,2,...,L_{y}\}$, then $A<B$ if $|A|<|B|$ or if $|A|=|B|$ then the least element in $A\cup B-A\cap B$ is contained in $A$. With this ordering on the subsets, we define the Hamiltonian on the extended Hilbert space as
$\begin{split}H&\equiv Diag(H_{0},H_{1},...,H_{L_{y}},...,H_{i_{1}...i_{k}},...,H_{1...L_{y}}),\qquad\{i_{1},...,i_{k}\}\subset\{1,...,L_{y}\}\end{split}$ (129)
where the entrees in the diagonal are ordered with respect to the subscript that is in one to one correspondence with the subsets of $\{1,2,...,L_{y}\}$. Hamiltonian $H_{i_{1}...i_{k}}$ denote the Hamiltonian Eq. (126) with defects inserted at rows labelled by $i_{1}$,…,$i_{k}$. Namely, the twisted Hamiltonian $H_{i_{1}...i_{k}}$ is defined as the Hamiltonian $H_{plaq}$ but with the product $S^{x}_{j}$ (and $S^{x}_{j+1}$) in the second term given a phase $(-1)$ when $j\in\{i_{1},...,i_{k}\}$. The untwisted Hamiltonian $H_{0}$ is same as the Hamiltonian $H_{plaq}$ in Eq. (126) without any defects inserted. In total there are $2^{L_{y}}$ copies of the Hilbert space in the total Hilbert space and $2^{L_{y}}$ Hamiltonians in the diagonal matrix that represent the Hamiltonian in the total Hilbert space. The dimension of the extended Hilbert space $\mathcal{H}$ is $dim(\mathcal{H}_{0})2^{L_{y}}$. We define the subsystem fermion parity on the total Hilbert space
@@
with $2^{L_{y}}$ copies of the subsystem fermion parity operator $(-1)^{\hat{F}_{i}}$ on the diagonal at $i^{\text{th}}$ row. In the extended Hilbert space, the Hamiltonian $H$ has other $\mathbb{Z}_{2}$ symmetries which we denote by $\eta_{i}$. These are $\eta_{i}=Diag\left(a_{0},a_{1},...,a_{L_{y}},...,a_{i_{1}..i_{k}},...,a_{1...L_{y}}\right)$ with
@@
However, we cannot take the Pauli $Z$ and $Y$ to be diagonal since they wouldn’t commute with the projection. Hence, we take them to be anti-diagonal, which in turn commute with the projection,
$Z\equiv\begin{pmatrix}&&Z\\ &\iddots&\\ Z&&\end{pmatrix},\qquad Y\equiv\begin{pmatrix}&&Y\\ &\iddots&\\ Y&&\end{pmatrix}.$ (143)
With these definitions of Pauli operators, the Hamiltonian $H$ in Eq. (129) is
@@ A.2 Majorana hypercubic models
$\begin{split}H_{HC}&=-\sum_{\begin{subarray}{c}i_{1}\neq L_{x_{1}}\\ ,i_{2},...,i_{d}\end{subarray}}\prod_{j_{2},...,j_{d}}Z_{i_{1},i_{2}+j_{2},...,i_{d}+j_{d}}Z_{i_{1}+1,i_{2}+j_{2},...,i_{d}+j_{d}}\\ &\hskip 28.45274pt-\sum_{i_{2},...,i_{d}}\prod_{j_{2},...j_{d}=0}^{1}\prod_{i_{1}=1}^{L_{x_{1}}}X_{i_{1},i_{2}+j_{2},...,i_{d}+j_{d}}\prod_{j_{2},...,j_{d}}Z_{L_{x_{1}},i_{2}+j_{2},...,i_{d}+j_{d}}Z_{1,i_{2}+j_{2},...,i_{d}+j_{d}}-\lambda\sum_{i_{1},...,i_{d}}X_{i_{1},...,i_{d}}.\end{split}$ (149)
@@
This is the hypercubic Ising model Hamiltonian with defects inserted along $x_{1}$ direction. The second term in Eq. (149) contain the defect lines. To gauge the subsystem fermion parity, we sum over defect configurations in the $x_{1}$ direction. Define an extended Hilbert space
@@
$\begin{split}\mathcal{H}&\equiv\mathcal{H}_{0}\bigoplus_{\{i_{1},...,i_{k}\}\neq\phi}\mathcal{H}_{i_{1}...i_{k}},\qquad i_{j}\in\{1,...,L_{x_{2}}...L_{x_{d}}\},\quad\{i_{1},...,i_{k}\}\subset\{1,...,L_{x_{2}}...L_{x_{d}}\}.\end{split}$ (150)
@@
We choose an ordering on the subsystem fermion parity lines in the $x_{1}$ direction. Any line in the $x_{1}$ direction is specified by the $d-1$ of the remaining coordinates $x_{2}$,…,$x_{d}$. Then we choose an ordering on the remaining coordinates $(x_{2},...,x_{d})<(x_{2}^{\prime},...,x_{d}^{\prime})$ if $x_{i}<x_{i}^{\prime}$ provided $x_{j}=x_{j}^{\prime}$ for all $j<i$, i.e. , ordering is chosen based on the first coordinate (reading from left) for which the two tuples disagree. Then these lines enumerated from 1 to $L_{x_{2}}...L_{x_{d}}$. Similar to the two dimensional case we define a total ordering on the subsets of $\{1,...,L_{x_{2}}...L_{x_{d}}\}$ with the definition of the total ordering exactly as in the case of two dimension. With respect to this ordering, we define the Hamiltonian on the extended Hilbert space
@@
$\begin{split}H\equiv Diag\left(H_{0},H_{1},...,H_{i_{1}...i_{k}},...,H_{1...(L_{x_{2}}...L_{x_{d}})}\right),\quad i_{j}\in\{1,...,L_{x_{2}}...L_{x_{d}}\},\quad\{i_{1},...,i_{k}\}\subset\{1,...,L_{x_{2}}...L_{x_{d}}\},\end{split}$ (151)
@@
where the entrees in the diagonal are ordered with respect to the subscript that is in one to one correspondence with the subsets of $\{1,2,...,L_{x_{2}}...L_{x_{d}}\}$. Hamiltonian $H_{i_{1}...i_{k}}$ denotes the Hamiltonian Eq. (149) with defects inserted on lines labelled by $i_{1}$,$i_{2}$,… up to $i_{k}$. Namely, the twisted Hamiltonian $H_{i_{1}...i_{k}}$ is defined as the Hamiltonian $H_{HC}$ but with the subsystem fermion parity lines enumerated as $j$ (according to our definition of enumeration) in the second term given a phase $(-1)$ when $j\in\{i_{1},...,i_{k}\}$. The untwisted Hamiltonian $H_{0}$ is same as the Hamiltonian $H_{plaq}$ in Eq. (126) without any defects inserted. In total, there are $2^{L_{x_{2}}...L_{x_{d}}}$ copies of the Hilbert space in the total Hilbert space. The rest of the analysis is a straightforward extension of the one we performed for the Majorana plaquette model in the previous subsection with $i$ taking values in $\{1,...,L_{x_{2}}...L_{x_{d}}\}$. After the projections Eq. (132), we obtain the total Hamiltonian Eq. (151) with the new Pauli $X$ and $Z$ operators similar as in Eqs. (136),(143)
@@ B.1 One dimension
Let us consider the operator
@@
$\begin{split}\overline{\mathbf{D}}_{even}&\equiv\left(\prod_{k=1}^{N-1}e^{i\frac{\pi}{4}Z_{2k-1}X_{2k}Z_{2k+1}}e^{i\frac{\pi}{4}Z_{2k}Z_{2k+2}}\right)e^{i\frac{\pi}{4}Z_{2N-1}X_{2N}Z_{1}}\frac{\left(1+\eta_{even}\right)}{2},\end{split}$ (155a)
$\begin{split}\overline{\mathbf{D}}_{odd}&\equiv\left(\prod_{k=1}^{N-1}e^{i\frac{\pi}{4}Z_{2k-2}X_{2k-1}Z_{2k}}e^{i\frac{\pi}{4}Z_{2k-1}Z_{2k+1}}\right)e^{i\frac{\pi}{4}Z_{2N-2}X_{2N-1}Z_{2N}}\frac{\left(1+\eta_{odd}\right)}{2}.\end{split}$ (155b)
Note that the pattern in the product of operators in Eq. (B.1) take an interesting form; alternating terms of the form $ZXZ$ and $ZZ$ in the exponents. In fact this pattern give rise to the desired Kennedy-Tasaki transformation. Explicitly, the action of our $\overline{\text{KT}}$ transformation is given by
@@
$\overline{\text{KT}}Z_{2i+1}$ $=Z_{2N}Y_{1}\prod_{k=1}^{i}X_{2k+1}Z_{2i+2}\overline{\text{KT}}^{\prime},$ (157a)
@@
$\mathbf{D}_{odd}^{\prime}$ $\equiv\Big{(}\prod_{k=1}^{N-1}e^{i\frac{\pi}{4}X_{2k-1}}e^{i\frac{\pi}{4}Z_{2k-1}Z_{2k+1}}\Big{)}e^{i\frac{\pi}{4}X_{2N-1}}\frac{(1-\eta_{odd})}{2},$ (159a)
$\mathbf{D}_{even}^{\prime}$ $\equiv\Big{(}\prod_{k=1}^{N-1}e^{i\frac{\pi}{4}X_{2k}}e^{i\frac{\pi}{4}Z_{2k}Z_{2k+2}}\Big{)}e^{i\frac{\pi}{4}X_{2N}}\frac{(1-\eta_{even})}{2}.$ (159b)
@@
$\overline{\text{KT}}^{\dagger}Z_{2i}=-Z_{2i-1}\prod_{k=i}^{N-1}X_{2k}Y_{2N}Z_{1}(\overline{\text{KT}}^{\prime\prime})^{\dagger},$ (160a)
$\overline{\text{KT}}^{\dagger}Z_{2i+1}=-Z_{2i}\prod_{k=i}^{N-2}X_{2k+1}Y_{2N-1}Z_{2N}(\overline{\text{KT}}^{\prime})^{\dagger}.$ (160b)
@@
$\overline{\text{KT}}^{2}=\mathbf{T}_{2}\frac{(1+\eta_{even})}{2}\frac{(1+\eta_{odd})}{2},$ (161)

### v2

@@ Acknowledgements.
The authors would like to thank Shu-Heng Shao, Yunqin Zheng and Mikhail Litvinov for useful discussions. APM and HS would like to thank Takuya Okuda for discussions on a project that uses similar ideas. This work was supported by the National Science Foundation under Award No. PHY 2310614. T.-C.W. also acknowledges the support by Stony Brook University’s Center for Distributed Quantum Processing.
@@ A.1 Majorana plaquette model
Hence, in total, there are $2^{L_{x}+L_{y}-1}$ subsystem symmetries, which in turn agree with the number of subsystem symmetries for the plaquette Ising model. Now we discuss gauging the subsystem fermion parity. We consider the Jordan-Wigner transformation for subsystem symmetric fermionic models in 2D defined in Ref. [57, 33],
@@
$\begin{split}\tilde{H}_{plaq}&=-\sum_{i\neq L_{x},j}Z_{i,j}Z_{i+1,j}Z_{i,j+1}Z_{i+1,j+1}-\sum_{j=1}^{L_{y}}\left(\prod_{i=1}^{L_{x}}X_{i,j}\right)\left(\prod_{i=1}^{L_{x}}X_{i,j+1}\right)Z_{L_{x},j}Z_{1,j}Z_{L_{x},j+1}Z_{1,j+1}-\lambda\sum_{i,j}X_{i,j}.\end{split}$ (128)
This is the plaquette Ising model Hamiltonian with defects along the horizontal direction. The second term in Eq. (128) contains defects inserted at two consecutive horizontal lines. Note that the subsystem fermion parity maps to the subsystem line symmetry of the plaquette Ising model under the Jordan-Wigner transformation Eq. (127),
@@
We perform a procedure similar to the procedure of gauging fermion parity in the free Majorana fermion model to obtain the transverse field Ising model in 1D [18]. First, let us define an extended Hilbert space.
$\mathcal{H}\equiv\mathcal{H}_{0}\bigoplus_{\{i_{1},...,i_{k}\}\neq\emptyset}\mathcal{H}_{i_{1}...i_{k}},\qquad\{i_{1},...,i_{k}\}\subset\{1,...,L_{y}\}$ (130)
where $\emptyset$ is the empty set and $i_{1}<...<i_{k}$. Now let us introduce a total ordering on the subsets of the set $\{1,2,...,L_{y}\}$. Suppose $A$ and $B$ are two subsets of $\{1,2,...,L_{y}\}$, then $A<B$ if $|A|<|B|$ ($|S|$ denote the cardinality of the set $S$) or if $|A|=|B|$ then the least element in $A\cup B-A\cap B$ is contained in $A$. With this ordering on the subsets, we define the Hamiltonian on the extended Hilbert space as
$\begin{split}H&\equiv Diag(H_{0},H_{1},...,H_{L_{y}},...,H_{i_{1}...i_{k}},...,H_{1...L_{y}}),\\ &\qquad\{i_{1},...,i_{k}\}\subset\{1,...,L_{y}\}\end{split}$ (131)
where the entrees in the diagonal are ordered with respect to the subscript that is in one-to-one correspondence with the subsets of $\{1,2,...,L_{y}\}$. Hamiltonian $H_{i_{1}...i_{k}}$ denote the Hamiltonian Eq. (128) with defects inserted at rows labelled by $i_{1}$,…,$i_{k}$. Namely, the twisted Hamiltonian $H_{i_{1}...i_{k}}$ is defined as the Hamiltonian $\tilde{H}_{plaq}$ but with the product $S^{x}_{j}$ ( or $S^{x}_{j+1}$) in the second term given a phase $(-1)$ when $j\in\{i_{1},...,i_{k}\}$( or $j+1\in\{i_{1},...,i_{k}\}$). The untwisted Hamiltonian $H_{0}$ is the same as the Hamiltonian $\tilde{H}_{plaq}$ in Eq. (128) without any defects inserted. In total there are $2^{L_{y}}$ isomorphic copies of the Hilbert space $\mathcal{H}_{0}$ in the total Hilbert space $\mathcal{H}$ and $2^{L_{y}}$ Hamiltonians in the diagonal matrix that represent the Hamiltonian $H$ in the total Hilbert space $\mathcal{H}$. The dimension of the extended Hilbert space $\mathcal{H}$ is $dim(\mathcal{H}_{0})2^{L_{y}}$. We define the subsystem fermion parity on the total Hilbert space
@@
with $2^{L_{y}}$ copies of the subsystem fermion parity operator $(-1)^{\hat{F}_{i}}$ at $i^{\text{th}}$ row on the diagonal. In the extended Hilbert space $\mathcal{H}$, the Hamiltonian $H$ has other $\mathbb{Z}_{2}$ symmetries which we denote by $\eta_{i}$. These are $\eta_{i}=Diag\left(a_{0},a_{1},...,a_{L_{y}},...,a_{i_{1}..i_{k}},...,a_{1...L_{y}}\right)$ with
@@
However, we cannot take the Pauli $Z$ and $Y$ to be diagonal since they wouldn’t commute with the projection. We define the new Pauli $Z$ and $Y$ operators that commute with the projection in terms of their matrix elements. This definition depend on the vertical position of the operators on the square lattice. Let $S\subset\{1,...,L_{y}\}$
$\left(Z_{-,j}\right)_{S\cup\{j\},S}=\left(Z_{-,j}\right)_{S,S\cup\{j\}}=Z_{-,j}\,,$ (136a)
$\left(Y_{-,j}\right)_{S\cup\{j\},S}=\left(Y_{-,j}\right)_{S,S\cup\{j\}}=Y_{-,j}\,,$ (136b)
where the index of subsets is numbered, according to the total ordering we defined above, starting from one. The Pauli $Z$ and $Y$ operators on the R.H.S. is acting on the Hilbert space $\mathcal{H}_{S}$ or $\mathcal{H}_{S\cup\{j\}}$. It is an easy exercise to see that above defined operators indeed satisfy the Pauli algebra. With these definitions of Pauli operators, the Hamiltonian $H$ in Eq. (131) is
@@ A.2 Majorana hypercubic models
$\begin{split}\tilde{H}_{HC}&=-\sum_{\begin{subarray}{c}i_{1}\neq L_{x_{1}}\\ ,i_{2},...,i_{d}\end{subarray}}\prod_{j_{2},...,j_{d}=0}^{1}Z_{i_{1},i_{2}+j_{2},...,i_{d}+j_{d}}Z_{i_{1}+1,i_{2}+j_{2},...,i_{d}+j_{d}}\\ &\hskip 28.45274pt-\sum_{i_{2},...,i_{d}}\prod_{j_{2},...j_{d}=0}^{1}\prod_{i_{1}=1}^{L_{x_{1}}}X_{i_{1},i_{2}+j_{2},...,i_{d}+j_{d}}\prod_{j_{2},...,j_{d}=0}^{1}Z_{L_{x_{1}},i_{2}+j_{2},...,i_{d}+j_{d}}Z_{1,i_{2}+j_{2},...,i_{d}+j_{d}}-\lambda\sum_{i_{1},...,i_{d}}X_{i_{1},...,i_{d}}.\end{split}$ (142)
@@
This is the hypercubic Ising model Hamiltonian with defects inserted along $x_{1}$ direction. The second term in Eq. (142) contains the defect lines. To gauge the subsystem fermion parity, we sum over defect configurations in the $x_{1}$ direction. Define an extended Hilbert space
@@
$\begin{split}\mathcal{H}&\equiv\mathcal{H}_{0}\bigoplus_{\{i_{1},...,i_{k}\}\neq\emptyset}\mathcal{H}_{i_{1}...i_{k}},\qquad i_{j}\in\{1,...,L_{x_{2}}...L_{x_{d}}\},\\ &\qquad\{i_{1},...,i_{k}\}\subset\{1,...,L_{x_{2}}...L_{x_{d}}\}.\end{split}$ (143)
@@
We choose an ordering on the subsystem fermion parity lines in the $x_{1}$ direction. Any line in the $x_{1}$ direction is specified by the $d-1$ of the remaining coordinates $x_{2}$,…,$x_{d}$. Then we choose an ordering on the remaining coordinates $(x_{2},...,x_{d})<(x_{2}^{\prime},...,x_{d}^{\prime})$ if $x_{i}<x_{i}^{\prime}$ provided $x_{j}=x_{j}^{\prime}$ for all $j<i$, i.e., ordering is chosen based on the first coordinate (reading from left) for which the two tuples disagree. Then these lines are enumerated from 1 to $L_{x_{2}}...L_{x_{d}}$. Similar to the two-dimensional case we define a total ordering on the subsets of $\{1,...,L_{x_{2}}...L_{x_{d}}\}$ with the definition of the total ordering exactly as in the case of two dimensions. With respect to this ordering, we define the Hamiltonian on the extended Hilbert space
@@
$\begin{split}H&\equiv Diag\left(H_{0},H_{1},...,H_{i_{1}...i_{k}},...,H_{1...(L_{x_{2}}...L_{x_{d}})}\right),\\ &\qquad\qquad\quad i_{j}\in\{1,...,L_{x_{2}}...L_{x_{d}}\},\\ &\qquad\quad\{i_{1},...,i_{k}\}\subset\{1,...,L_{x_{2}}...L_{x_{d}}\},\end{split}$ (144)
@@
where the entrees in the diagonal are ordered with respect to the subscript that is in one-to-one correspondence with the subsets of $\{1,2,...,L_{x_{2}}...L_{x_{d}}\}$. Hamiltonian $H_{i_{1}...i_{k}}$ denotes the Hamiltonian Eq. (142) with defects inserted on lines labelled by $i_{1}$,$i_{2}$,… up to $i_{k}$. Namely, the twisted Hamiltonian $H_{i_{1}...i_{k}}$ is defined as the Hamiltonian $\tilde{H}_{HC}$ but with the subsystem fermion parity lines enumerated as $j$ (according to our definition of enumeration) in the second term given a phase $(-1)$ when $j\in\{i_{1},...,i_{k}\}$. The untwisted Hamiltonian $H_{0}$ is the same as the Hamiltonian $\tilde{H}_{HC}$ in Eq. (128) without any defects inserted. In total, there are $2^{L_{x_{2}}...L_{x_{d}}}$ isomorphic copies of the Hilbert space $\mathcal{H}_{0}$ in the total Hilbert space $\mathcal{H}$. The rest of the analysis is a straightforward extension of the one we performed for the Majorana plaquette model in the previous subsection with $i$ taking values in $\{1,...,L_{x_{2}}...L_{x_{d}}\}$. After the projections Eq. (134), we obtain the total Hamiltonian Eq. (144) with the new Pauli $X$ and $Z$ operators similar as in Eqs. (135),(136)
@@ B.1 One dimension
We consider a ring with $2M$ sites. Let us consider the operator
@@
$\begin{split}\overline{\mathbf{D}}_{even}&\equiv\left(\prod_{k=1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-1}e^{i\frac{\pi}{4}Z_{2k-1}X_{2k}Z_{2k+1}}e^{i\frac{\pi}{4}Z_{2k}Z_{2k+2}}\right)\\ &\quad\times e^{i\frac{\pi}{4}Z_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-1}X_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}Z_{1}}\frac{\left(1+\eta_{even}\right)}{2},\end{split}$ (148a)
$\begin{split}\overline{\mathbf{D}}_{odd}&\equiv\left(\prod_{k=1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-1}e^{i\frac{\pi}{4}Z_{2k-2}X_{2k-1}Z_{2k}}e^{i\frac{\pi}{4}Z_{2k-1}Z_{2k+1}}\right)\\ &\times e^{i\frac{\pi}{4}Z_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-2}X_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-1}Z_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}}\frac{\left(1+\eta_{odd}\right)}{2}.\end{split}$ (148b)
Note that the pattern in the product of operators in Eq. (148) takes an interesting form; alternating terms of the form $ZXZ$ and $ZZ$ in the exponents. In fact, this pattern gives rise to the desired Kennedy-Tasaki transformation. Explicitly, the action of our $\overline{\text{KT}}$ transformation is given by
@@
$\overline{\text{KT}}Z_{2i+1}$ $=Z_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}Y_{1}\prod_{k=1}^{i}X_{2k+1}Z_{2i+2}\overline{\text{KT}}^{\prime},$ (150a)
@@
$\mathbf{D}_{odd}^{\prime}$ $\equiv\Big{(}\prod_{k=1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-1}e^{i\frac{\pi}{4}X_{2k-1}}e^{i\frac{\pi}{4}Z_{2k-1}Z_{2k+1}}\Big{)}e^{i\frac{\pi}{4}X_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-1}}\frac{(1-\eta_{odd})}{2},$ (152a)
$\mathbf{D}_{even}^{\prime}$ $\equiv\Big{(}\prod_{k=1}^{N-1}e^{i\frac{\pi}{4}X_{2k}}e^{i\frac{\pi}{4}Z_{2k}Z_{2k+2}}\Big{)}e^{i\frac{\pi}{4}X_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}}\frac{(1-\eta_{even})}{2}.$ (152b)
@@
$\overline{\text{KT}}^{\dagger}Z_{2i}=-Z_{2i-1}\prod_{k=i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-1}X_{2k}Y_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}Z_{1}(\overline{\text{KT}}^{\prime\prime})^{\dagger},$ (153a)
$\overline{\text{KT}}^{\dagger}Z_{2i+1}=-Z_{2i}\prod_{k=i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-2}X_{2k+1}Y_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}-1}Z_{2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}(\overline{\text{KT}}^{\prime})^{\dagger}.$ (153b)
@@
$\overline{\text{KT}}^{2}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}e^{i\pi M}}\mathbf{T}_{2}\frac{(1+\eta_{even})}{2}\frac{(1+\eta_{odd})}{2},$ (154)